OpenStudy (anonymous):

(16x^2 -25) divided by (4x + 5)

OpenStudy (anonymous):

\[\frac{ 16x^{2} - 25 }{ 4x+5 } = \frac{ (4x+5)\times(4x-5)) }{ 4x+5 } = 4x-5\]

OpenStudy (anonymous):

thank you!

OpenStudy (dawnr):

I was about to write that :D

OpenStudy (anonymous):

Don't worry about it!